[Brahms-l] Shift report 20060613 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue Jun 13 2006 - 00:04:17 EDT
Shifter leader was dp
Shifters were CN


We have beam for 45 minutes. 
We stopped taking data because the background was very big. afterthat beam was dumped and  change our FS polarity. Now all magnets have B polarity. 
Afterthat there was no real beam. PHENIX test their spin rotator settings. 
We took non tpc pedestal and tpc pedestal runs.

We have archived all the supermon plots. They can be found on the topshelf near the fire exit doors in the counting house

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Jun 13 00:04:48 2006

This archive was generated by hypermail 2.1.8 : Tue Jun 13 2006 - 00:05:02 EDT